OpenStudy (allieatthedisco):

The coordinates of point T are given. The midpoint of line ST is (5,-8). Find the coordinates of point S where T is (0,4)

OpenStudy (ghazi):

(x+0)/2=5 and (y-8)/2=4

OpenStudy (ghazi):

can you do this?

OpenStudy (allieatthedisco):

i think so, give me a second

OpenStudy (ghazi):

what you got?

OpenStudy (allieatthedisco):

ok i have no clue what to do. is there any other way you could explain it?

OpenStudy (ghazi):

well that was just mean of two numbers

OpenStudy (ghazi):

\[\frac{ x+0 }{ 2 }=5 and \frac{ y-8 }{ 2 }=4\]

OpenStudy (ghazi):

x=10 and y=8+8=16

OpenStudy (allieatthedisco):

how do you find what x and y equal though? that's where im getting lost

OpenStudy (ghazi):

multiply both sides by 2 for x and for y

OpenStudy (allieatthedisco):

oh ok, thank you